40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 9009     Entry time: Tue Aug 13 21:49:32 2013
Author: Koji 
Type: Summary 
Category: General 
Subject: Testing new AG4395A network analyzer 

New AG4395, sn MY41101114  for West Bridge Labs was delivered. For the test purpose it is at the 40m now.

I made a series of tests in order to find anything broken.

Network analyzer test

- RF out / Rch test

RF out directly connected to R input channel.
The received power at the R-ch was measured while the output was swept from 10Hz to 500MHz.

The RF power was changed from -50dBm to +15dBm with +10dBm increment (but the last one).

The attenuator setting was changed from 50dB to 0dB.

=> The configured output power was properly detected by the R channel.

=> RF output is producing the signal properly. R-ch is detecting the produced signal properly.

- Ach/Bch test

Same test as above for Ach and Bch 

=> Same result as above

=> A-ch and B-ch are detecting the produced signal properly.

- Transfer function test

Connect a power splitter to the RF out. Detect the split signals by R-ch and A-ch

=> Measurement is at around 0dB +/- 1dB up to 500MHz.

Same measurement for B-ch

=> Same result

=> A/R and B/R indicates proper transfer function measurements.

- Calibration

RF out was split in to two. One was connected to R-ch. The other was connected to A-ch.
The thru response calibration was run.

=> The thru calibration was performed properly. 

- Practical tranfer function measurements.

In the above calibration setup, various RF filters were inserted in the Ach path.

The measured data was extracted via GPIB connection.

=> Practical transfer function measurements were performed.

=> GPIB connectivity was confirmed

 

External reference test

- External 10MHz reference from an SRS frequency counter was connected to Ext Ref In

=> Ext Ref indicator on the screen appeard

=> The internal oscillator seemed to be locked to the external reference in

 

 

Spectrum analyzer test

- Measured the signals from DS345 by R/A/B ch

Sinusoidal signal (1V) swept from 10MHz to 30Mhz

=> Corresponding moving peak was detected in each case

- Noise level measurement

R/A/B channels were terminated. The attenuation at each port was set to 0dB.

Frequency span was changed between 500MHz, 10MHz, 100kHz, 1kHz.

=> Noise level of ~10nV/rtHz between 0.1-500MHz was confirmed. All R/A/B channels have the same performance.

Attachment 1: AG4395A_noise.pdf  80 kB  Uploaded Tue Aug 13 23:02:07 2013  | Hide | Hide all
AG4395A_noise.pdf
ELOG V3.1.3-